Question 1 (45%): In a perfectly competitive market, the demand curve is given as: Q=100-5P, the supply curve is given as Q=3P-12.

I. Compute the total social surplus of this market. (10%)

II. If the government impose a tax on the producers, and the tax rate is $2 per unit produced. What is the deadweight loss? (10%)

III. If the government impose a tax on the consumers, and the tax rate is $2 per unit purchased, graphically show the change in the market equilibrium and the deadweight loss. (15%)

IV. If the government provide a subsidy of $4 per unit to the producers, what is the deadweight loss? (10%)
Answer #1

demand: Q=100-SP Supply : 8 3P-12 Total Surkhes - Consumer Producer Susplus t Susphes. 4 V 160 30 Equilibrium print 100-5P =

Date PONO Producer Surplus - 1.x (14.4) 430. - 1 x 10 x 30 15 = 150 Social Susflus = 90+150 = 240. 61 Tax is imposed on boo d

Q = 14 x 15 14X IS 8 26.95 = .1 x lax & BQuy 21x2x ( 30-26.25) | DWL = 3.75 DWL-

D ca (0) Deadweight loss 26.55 30 100 & demand curve 100 - SCP+2) = 100 -5P - 10 10 - 90 - SP y cume = Q = 3P-12 90-5P = 3P-1

Date 20 Supply. 9 - 3P+4)-12 Q = 3P + 12-12 19 = 32 Dem and : Q-100-SP In equilibrium Qs = 0 Q = 3p/ = 1300-se 3P = 8P = P =